OpenStudy (anonymous):

(5x+6)=(7x-12) Solve???

OpenStudy (anonymous):

Simplifying (5x + 6) = (7x + -12) Reorder the terms: (6 + 5x) = (7x + -12) Remove parenthesis around (6 + 5x) 6 + 5x = (7x + -12) Reorder the terms: 6 + 5x = (-12 + 7x) Remove parenthesis around (-12 + 7x) 6 + 5x = -12 + 7x Solving 6 + 5x = -12 + 7x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-7x' to each side of the equation. 6 + 5x + -7x = -12 + 7x + -7x Combine like terms: 5x + -7x = -2x 6 + -2x = -12 + 7x + -7x Combine like terms: 7x + -7x = 0 6 + -2x = -12 + 0 6 + -2x = -12 Add '-6' to each side of the equation. 6 + -6 + -2x = -12 + -6 Combine like terms: 6 + -6 = 0 0 + -2x = -12 + -6 -2x = -12 + -6 Combine like terms: -12 + -6 = -18 -2x = -18 Divide each side by '-2'. x = 9 Simplifying x = 9

OpenStudy (anonymous):

69

OpenStudy (anonymous):

so my answer would be 51.

OpenStudy (anonymous):

This is how the problem is shown.

OpenStudy (anonymous):

yes its 51